# Inference for Regression with

### 23 Feb 2024 v1

@@ 3.2.1 Fixed Population
$\hat{\bm{\gamma}}\to_{p}\left(\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]+\mathbb{E}\left[\frac{1}{C_{i}}\right]\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\right)^{-1}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\,\bm{\gamma}\,.$
In particular, if $\mathbb{E}[C_{i}^{-1}]$ is small, then
$\hat{\bm{\gamma}}\to_{p}\bm{\gamma}-\mathbb{E}\left[\frac{1}{C_{i}}\right]\left(\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbf{I}\right)\bm{\gamma}+O\left(\mathbb{E}\left[\frac{1}{C_{i}}\right]^{2}\right)\,.$
The first part of Theorem 1 shows that the measurement error introduced by regressing $Y_{i}$ on $\hat{\bm{\theta}}_{i}$ instead of the true (infeasible) $\bm{\theta}_{i}$ makes the estimator $\hat{\bm{\gamma}}$ inconsistent. How well we can impute the true latent $\bm{\theta}_{i}$ for each observation $i$ depends on the amount of unstructured data $C_{i}$. Because each $C_{i}$ is finite, each $\hat{\bm{\theta}}_{i}$ has a measurement error that doesn’t disappear when the number of observations $n$ becomes large. As a consequence, $\hat{\bm{\gamma}}$ is biased (even asymptotically).
More constructively, Theorem 1 shows that what is important for controlling bias is not the average amount of unstructured data $\mathbb{E}[C_{i}]$ but rather the mean reciprocal amount $\mathbb{E}[C_{i}^{-1}]$. This makes intuitive sense, as the measurement error in $\hat{\bm{\theta}}_{i}$ decays with $C_{i}$ but the rate of decay decreases with $C_{i}$. If the population contains a larger share of observations with small $C_{i}$, then larger measurement errors in $\hat{\bm{\theta}}_{i}$ will be more prevalent and $\hat{\bm{\gamma}}$ will have a larger bias [^12] To put it differently, increasing the amount of unstructured data for the observations with small $C_{i}$ will have a larger effect on the expected average size of the measurement error than if the additional data was collected for the observations with large $C_{i}$. Consequently, the across-observation distribution of $C_{i}$ matters beyond its mean. It is important to emphasize that even if most observations have a large $C_{i}$ but a small mass do not (meaning that $\mathbb{E}[C_{i}^{-1}]$ may still be large), then the noise in $\hat{\bm{\theta}}_{i}$ from the observations with small $C_{i}$ can still substantially bias $\hat{\bm{\gamma}}$.
The second part of Theorem 1 shows that when all observations have a large amount of unstructured data (so that $\mathbb{E}[C_{i}^{-1}]$ is small), the first-order effect is a bias of size
$-\mathbb{E}\left[\frac{1}{C_{i}}\right]\,\left(\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbf{I}\right)\bm{\gamma}.$
Thus, the first-order effect of bias is proportional to $\mathbb{E}[C_{i}^{-1}]$.
@@ 3.2.2 Sequence of Populations
We now build on this insight to consider a sequence of populations where the amount of unstructured data per observation becomes larger as the sample size $n$ increases. This asymptotic framework is designed to shed light on how $\hat{\bm{\gamma}}$ behaves when there is a relatively large number of observations and there is a large amount of unstructured data per observation. In this scenario, the measurement errors for each observation are small but their cumulative effect may not be completely ignorable relative to sampling error.
Formally, we consider a sequence of populations indexed by sample size $n$. In each population, we keep the distribution of $(Y_{i},\bm{\theta}_{i},\mathbf{q}_{i})$ conditional on $C_{i}$ fixed and as described in Section 3.1. We also maintain the assumption that $\mathbf{x}_{i}$ is given by the topic model (5). However, we let the marginal distribution of $C_{i}$ to change with the sample size $n$ to allow the amount of unstructured data per observation to become large as the sample size $n$ increases. Specifically, we consider a framework in which
@@
The case $\kappa=0$ corresponds to a setting in which the amount of unstructured data per observation is of much larger order than sample size. Consequently, measurement error is of smaller order (asymptotically) than sampling error. In this case, our theory implies that the two-step strategy leads to valid inference. That is, the measurement error introduced by regressing $Y_{i}$ on $\hat{\bm{\theta}}_{i}$ instead of $\bm{\theta}_{i}$ can effectively be ignored and standard inference can proceed treating the $\hat{\bm{\theta}}_{i}$ as if they are the true $\bm{\theta}_{i}$.
The case $\kappa\in(0,\infty)$ is the critical case in which there is a large, but not overwhelming, amount of unstructured data per observation. This case mimics many empirically realistic designs where measurement error and sampling error are both small but non-negligible. We show in this case that $\hat{\bm{\gamma}}$ is consistent but standard two-step inference is invalid. In particular, the asymptotic distribution of $\hat{\bm{\gamma}}$ has the correct variance but its center is shifted due to measurement error bias. Consequently, confidence intervals based on the usual two-step strategy have the correct width but incorrect centering, and therefore have a coverage rate that is smaller than nominal coverage. [^13] The case $\kappa=+\infty$ corresponds to a setting where measurement error is of larger order than sampling error. Here $\hat{\bm{\gamma}}$ is consistent provided $\mathbb{E}[C_{i}^{-1}]\to 0$ but two-step inference is invalid because bias is of larger order than sampling uncertainty. In that case, the coverage rates of standard OLS confidence intervals asymptote to zero as the sample size $n$ becomes large.
@@
    $\mathbf{B}$, $\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]$, and $\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]$ have full rank.
@@
    $\mathbb{E}\left[|\varepsilon_{i}|^{2+\delta}\right]<\infty$ for some $\delta>0$.
@@
Assumption 2(i) is mostly the same as Assumption 1(i) except we also require that $\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]$ has full rank so that the asymptotic variance of $\hat{\bm{\gamma}}$ is well defined. Parts (ii) and (iii) strengthen parts (ii) and (iii) of Assumption 1 to require convergence a faster-than-root-$n$ rate. This is really to simplify the derivation: if convergence occurs instead at a root-$n$ rate, then additional terms may distort the asymptotic distribution further. Nevertheless we believe these two conditions are broadly satisfied. For instance, part (ii) is mild in view of known convergence rates established for estimators of $\mathbf{B}$ [^14] Bing et al., (2020), Wu et al., (2023), and Ke and Wang, (2022) derive finite-sample upper bounds for various estimators $\hat{\mathbf{B}}$ of $\mathbf{B}$. Each of their results implies the corresponding estimator $\hat{\mathbf{B}}$ converges at the optimal rate $(nC)^{-1/2}$ (up to log terms) where, for simplicity, the $C_{i}$ are all of the same order $C$. Hence, all estimators $\hat{\mathbf{B}}$ converge faster than $n^{-1/2}$ when $C$ grows with $n$, as we have here by (7). As before, part (iii) is made to simplify the derivation but is also not vacuous: given any estimator $\hat{\mathbf{B}}$ proposed in the literature satisfying part (ii), one could construct $\hat{\bm{\theta}}_{i}$ directly by setting $\hat{\bm{\theta}}_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}$, in which case $\sqrt{n}\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|=0$. As before, the first condition in (i) and parts (ii) and (iii) hold trivially for the pure multinomial model because $\hat{\mathbf{B}}=\mathbf{B}=\mathbf{I}$ and $\hat{\mathbf{p}}_{i}=\hat{\bm{\theta}}_{i}$. Part (iv) is standard for inference for regression under conditional heteroskedasticity (e.g., White, (1980)). Finally, Part (v) is made to simplify technical derivations and can be relaxed. This assumption means that for each $n$, $C_{i}$ is supported on $[c(\log n)^{1+\epsilon},\infty)$ for some constants $c,\epsilon>0$. This condition is much weaker than the conventional assumption that all $C_{i}$ grow at the same rate $C$ (Bing et al., 2020, Wu et al., 2023, Ke and Wang, 2022) which, in view of (7), would imply that $C_{i}$ is supported on $[cn^{1/2},\infty)$. This condition is only used to establish consistency of Eicker–Huber–White standard errors and is not required for asymptotic normality.
Our second main result shows that the OLS estimator of $\bm{\gamma}$ in equation (6) is consistent and derives its asymptotic distribution.
@@
$\sqrt{n}\left(\hat{\bm{\gamma}}-\bm{\gamma}\right)+\kappa\left(\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbf{I}\right)\bm{\gamma}\\[4.0pt] \to_{d}N\left(\bm{0},\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}\right)$ (8)
@@
$\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)^{-1}\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1},$ (9)
where $\hat{\varepsilon}_{i}=Y_{i}-\hat{\bm{\theta}}_{i}^{T}\hat{\bm{\gamma}}$.
Theorem 2 shows that inference is valid when $\kappa=0$. In this case, we have
$\sqrt{n}\left(\hat{\bm{\gamma}}-\bm{\gamma}\right)\to_{d}N\left(\bm{0},\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]^{-1}\right).$
The OLS estimator obtained by regressing $Y_{i}$ on $\hat{\bm{\theta}}_{i}$ therefore has the same asymptotic distribution as the (infeasible) OLS estimator obtained by regressing $Y_{i}$ on the true latent $\bm{\theta}_{i}$. The reason is that $\kappa=0$ corresponds to a scenario where measurement error is of smaller order than sampling error. Moreover, the usual Eicker–Huber–White standard errors computed using the estimates $\hat{\bm{\theta}}_{i}$ are consistent. Hence, measurement error can effectively be ignored when performing inference on $\bm{\gamma}$.

### 24 Mar 2024 v2

@@ 3.2.1 Fixed Population
$\hat{\bm{\psi}}\to_{p}\bigg{(}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]+\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left(\mathbf{Q}_{\mathbf{B}}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{Q}_{\mathbf{B}}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right]\bigg{)}^{-1}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\,\bm{\psi}\,.$
In particular,
$\mathrm{plim}(\hat{\bm{\psi}})\approx\bm{\psi}-\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{b}$
for $\mathbb{E}\left[C_{i}^{-1}\right]$ small, where
$\bm{b}=\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\left[\begin{array}[]{c}\bm{S}\left(\mathbf{Q}_{\mathbf{B}}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{Q}_{\mathbf{B}}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}\bm{\gamma}\\ \mathbf{0}\end{array}\right].$
Theorem 1 shows that $\hat{\bm{\psi}}$ is inconsistent due to the measurement error in $\hat{\bm{\theta}}_{i}$. The reason is that each $\hat{\bm{\theta}}_{i}$ has a measurement error that doesn’t disappear when the number of observations $n$ becomes large. As a consequence, $\hat{\bm{\psi}}$ is asymptotically biased.
More constructively, Theorem 1 shows that bias is proportional to the average inverse amount of unstructured data per observation, $\mathbb{E}\left[C_{i}^{-1}\right]$. Because this is an inverse relationship, increasing the amount of unstructured data for observations with small $C_{i}$ will reduce bias by more than if the additional data was collected for the observations with large $C_{i}$. Consequently, the across-observation distribution of $C_{i}$ matters beyond its mean. In particular, bias may be large if most observations have large $C_{i}$ but a small mass have small $C_{i}$.
@@ 3.2.2 Sequence of Populations
We now consider a sequence of populations where the amount of unstructured data per observation becomes larger as the sample size $n$ increases. This asymptotic framework is designed to shed light on how $\hat{\bm{\psi}}$ behaves in empirically realistic settings where there are a relatively large number of observations and there is a large amount of unstructured data per observation. In this scenario, the measurement errors for each observation are small but their cumulative effect may not be ignorable relative to sampling error.
Formally, we consider a sequence of populations indexed by sample size $n$. In each population, we keep the distribution of $(Y_{i},\bm{\theta}_{i},\mathbf{q}_{i})$ conditional on $C_{i}$ fixed and as described in Section 3.1. We also maintain the assumption that $\mathbf{x}_{i}$ is given by the topic model (5). However, we let the marginal distribution of $C_{i}$ change with the sample size $n$ to allow the amount of unstructured data per observation to become large as the sample size $n$ increases. Specifically, we consider a framework in which
@@
The case $\kappa=0$ corresponds to a setting in which the amount of unstructured data per observation is of much larger order than sample size. Consequently, measurement error is of smaller order than sampling error. In this case, our theory implies that the two-step strategy leads to valid inference. That is, the measurement error in $\hat{\bm{\theta}}_{i}$ can effectively be ignored and standard inference can proceed treating the $\hat{\bm{\theta}}_{i}$ as if they are the true $\bm{\theta}_{i}$.
The case $\kappa\in(0,\infty)$ is the critical case in which there is a large, but not overwhelming, amount of unstructured data per observation. This case mimics many empirically realistic designs where measurement error and sampling error are both small but non-negligible. We show in this case that $\hat{\bm{\psi}}$ is consistent but standard two-step inference is invalid. In particular, the asymptotic distribution of $\hat{\bm{\psi}}$ has the correct variance but its center is shifted due to measurement error bias. Consequently, confidence intervals based on the usual two-step strategy have the correct width but incorrect centering, and therefore have a coverage rate that is smaller than nominal coverage. [^10] The case $\kappa=+\infty$ corresponds to a setting where measurement error is of larger order than sampling error. Here $\hat{\bm{\psi}}$ is consistent provided $\mathbb{E}[C_{i}^{-1}]\to 0$ but two-step inference is invalid because bias is of larger order than sampling uncertainty. In that case, the coverage rates of standard OLS confidence intervals asymptote to zero as the sample size $n$ becomes large.
@@
    $\mathbf{B}$ has full rank.
@@
    $\mathbb{E}\left[\|\mathbf{q}_{i}\|^{4}\right]<\infty$ and $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ has full rank.
@@
Assumption 2(i) is the same as Assumption 1(i). Assumption 2(ii)-(iii) strengthens Assumption 1(ii)-(iii) to require convergence a faster-than-root-$n$ rate. We believe Assumption 2(ii) is broadly satisfied in view of known convergence rates for estimators of $\mathbf{B}$ [^11] Bing et al., (2020), Wu et al., (2023), and Ke and Wang, (2022) derive finite-sample guarantees for different estimators $\hat{\mathbf{B}}$ of $\mathbf{B}$. Each of their results implies the corresponding estimator $\hat{\mathbf{B}}$ converges at the optimal rate $(nC)^{-1/2}$ (up to log terms) where, for simplicity, the $C_{i}$ are all of the same order $C$. Hence, all estimators $\hat{\mathbf{B}}$ converge faster than $n^{-1/2}$ when $C$ grows with $n$, as we have here by (7). Assumption 2(iii) is made to simplify derivations but is not vacuous: given any estimator $\hat{\mathbf{B}}$, one could set $\hat{\bm{\theta}}_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}$, in which case $\sqrt{n}\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|=0$. As before, Assumption 2(i)-(iii) holds trivially for the pure multinomial model because $\hat{\mathbf{B}}=\mathbf{B}=\mathbf{I}$ and $\hat{\mathbf{p}}_{i}=\hat{\bm{\theta}}_{i}$. Assumptions 2(iv) and 2(v) are standard. Assumption 2(vi) is made to simplify technical derivations and can be relaxed. It implies that $C_{i}$ is supported on $[c(\log n)^{1+\epsilon},\infty)$ for some $c,\epsilon>0$. This is weaker than the conventional assumption that all $C_{i}$ grow at the same rate $C$ (Bing et al., 2020, Wu et al., 2023, Ke and Wang, 2022) which, in view of (7), would imply that $C_{i}$ is supported on $[cn^{1/2},\infty)$. This condition is only used to establish consistency of standard errors.
Our second main result shows that $\hat{\bm{\psi}}$ is consistent, derives its asymptotic distribution, and establishes consistency of standard errors. Recall the definition of $\bm{b}$ from Theorem 1. Let $\hat{\varepsilon}_{i}=Y_{i}-\hat{\bm{\xi}}_{i}^{T}\hat{\bm{\psi}}$.
@@
$\sqrt{n}(\hat{\bm{\psi}}-\bm{\psi})\\ \to_{d}N\left(\kappa\bm{b},\,\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\right),$ (8)
@@
$\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\varepsilon}_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}.$ (9)
Theorem 2 shows that two-step inference is valid when $\kappa=0$. In this case, we have
$\sqrt{n}(\hat{\bm{\psi}}-\bm{\psi})\to_{d}N\left(\bm{0},\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]^{-1}\right).$
Here $\hat{\bm{\psi}}$ has the same asymptotic distribution as the (infeasible) OLS estimator obtained by regressing $Y_{i}$ on the true latent $\bm{\theta}_{i}$. Moreover, Eicker–Huber–White standard errors computed using $\hat{\bm{\theta}}_{i}$ are consistent. The reason is that $\kappa=0$ corresponds to a scenario where measurement error is of smaller order than sampling error. Hence, measurement error can effectively be ignored when performing inference on $\bm{\psi}$.